0 0
Add a comment Improve this question Transcribed image text
Answer #1

Guass law :

The net flux crossing via any closed surface is equal to "net charge inclosed by that surface divided by permittivity of free space "
net charge inclosed by that surface =
Qin = q1+ 4 q2 =
+8 *1012 + (4*(- 2*10-12 ) C = 0C
so
net flux = 0